# library.bib

### v1

@@ Acknowledgements
This work was partly conducted during a sabbatical stay by GRF at the Global Research Centre for Diverse Intelligences at the University of St. Andrews, supported by PASPA-DGAPA grant from the Universidad Nacional Autónoma de México. RSW was supported by the EPSRC Centre for Doctoral Training in Mathematical Modelling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1). SSA was supported by a postdoctoral grant (Estancias posdoctorales por México) from the Secretaría de Ciencia, Humanidades, Tecnología e Innovación (SECIHTI). Fieldwork was carried out under research permits DGVS00910/13 and DGVS02716/14 from the Secretaría de Medio Ambiente y Recursos Naturales (SEMARNAT).
@@
Research Institute on Applied Mathematics and Systems, National Autonomous University of Mexico
@@
Institute of Physics, National Autonomous University of Mexico
@@
Correspondence to Gabriel Ramos-Fernandez

### v3

@@ Acknowledgements
This work was partly conducted during a sabbatical stay by GRF at the Global Research Centre for Diverse Intelligences at the University of St. Andrews, supported by a PASPA-DGAPA grant from the Universidad Nacional Autónoma de México. RSW was supported by the EPSRC Centre for Doctoral Training in Mathematical Modelling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1). SSA was supported by a postdoctoral grant (Estancias posdoctorales por México) from the Secretaría de Ciencia, Humanidades, Tecnología e Innovación (SECIHTI). Fieldwork was carried out under research permits DGVS00910/13 and DGVS02716/14 from the Secretaría de Medio Ambiente y Recursos Naturales (SEMARNAT). We thank Lev Guzman-Vargas, Alvaro Zabaleta-Ortega and two anonymous reviewers for their helpful comments on previous versions of the manuscript.
@@
Instituto de Investigaciones en Matemáticas Aplicadas y en Sistemas, Universidad Nacional Autónoma de México
@@
Instituto de Física, Universidad Nacional Autónoma de México
@@
Correspondence to Gabriel Ramos-Fernandez.